# OpenDebateEvidence: A Massive-Scale Argument Mining and Summarization Dataset

### v2

@@ A.1 Representation Bias
While the OpenDebateEvidence dataset is extensive, it may not fully represent the diversity of argumentation styles and topics across all debate communities. The dataset primarily includes evidence from American high school and college debates, and therefore might not capture the nuances of debates in other cultural or educational contexts, or in other languages.
@@ Appendix B Ethics Statement
The OpenDebateEvidence dataset, presented in this paper, derives from openly shared debate evidence across various educational forums and debate formats. This dataset strictly adheres to the principles of fair use, focusing on academic and research intent. The files which make up OpenDebateEvidence have been hosted online in some cases for over a decade without any known ethical issues arising as a result of it.
@@ Appendix C Social Impacts
In summary, we believe our work will accelerate research, improve model evaluation and training, and ultimately enhance the capabilities of language models in handling complex argumentative texts, with no foreseeable negative societal impacts.
@@ D.3 Orthogonalization
We used the notebook for this process.
@@ E.1.2 LLM as Judge Prompt
Score the abstract from 0-10 on it’s how well it supports the documents argument, and on its general quality
@@ E.2.2 LLM as Judge Prompt
SYSTEM PROMPT: You are a lawmaker.
USER PROMPT:
@@
Score the abstract from 0-10 on it’s how well it supports the documents argument, and on its general quality
@@ F.1.1 Dataset Summary
  • Massively increased size (about 25X the size of DebateSum), including nearly all debate evidence ever open sourced over the past 20 years from High School and College Public Forum, Policy, and Lincoln Douglas debate leagues
@@ F.1.2 Supported Tasks and Leaderboards
This dataset is useful for text generation, summarization, information retrieval, question answering, and related tasks. This dataset is further highly useful as a "trustworthy" dataset. All evidence within it has corresponding citations and is in general "factual" or grounded in facts. We do the evaluation in our paper, establishing the first "leaderboard" for measuring the performance of models trained on this dataset.
@@ F.1.3 Languages
English with very minor exceptions (i.e. evidence from performance cases using non-English evidence to make anti-colonialist arguments)
@@ F.1.6 Dataset Format
This dataset is originally contained in csv files, which were auto-converted into the parqueet dataset format by Huggingface. It’s available for download and consumption in both formats.
@@ F.1.8 Discussion of Biases
Competitive debate at the highest levels has increasingly rewarded teams who cite particular subfields of philosophy. A partial list of these highly represented topics is given below
@@
These cannons are dominated by so called "left-wing" thinkers and have mostly marginalized so called "right-wing" thinkers within them with some notable exceptions
@@ F.1.11 Personal Information
F.1.12 Licensing Information
All data within this dataset is clearly used within an extracurricular, educational activity. This means that any "copyright" issues from reproduction of copyrighted articles within the dataset are allowed and exempted under US copyright law. The OpenCaseList project fully blesses this work and has published this evidence with a permissive, MIT license.
F.1.13 Author Statement
We, the authors, bear all responsibility in case of violation of rights, etc., and we confirm that the data is licensed with an MIT license.
@@ F.1.13 Author Statement
tag Biased abstractive summary of the evidence / argument made by debater with evidence.
@@
pocket String indicating the virtual “pocket” (top level section, usually the speech name) in which the
@@
hat String indicating the virtual “hat” (medium level section, usually the broad type of argument)
@@
block String indicating the virtual “block” (low level section, usually the specific type of argument)
@@
Year (Top 5) 2020 850,607
2021 787,685
2019 609,703
2018 423,378
2017 258,742
Event cx 2,380,600
ld 1,164,132
pf 26,366
Caselist DisplayName (Top 3) HS LD 2020-21 383,489
HS LD 2021-22 326,166
HS Policy 2021-22 262,292
TeamSize 2 2,406,966
1 1,164,132
Level hs 2,144,757
college 1,426,341
State (Top 5) None 1,875,021
CA 450,734
TX 335,087
GA 108,011
IL 94,084
Side N 1,992,850
A 1,578,248
DuplicateCount (Top 5) 1 1,176,971
2 114,268
3 92,667
4 78,804
5 67,166

### v3

@@ 5 Conclusion
Acknowledgments
The authors want to thank and acknowledge the following individuals who have made this work possible:
  • Aaron Hardy - The long-time steward of the OpenCaseList project from which we gathered OpenDebateEvidence.
  • Rotem Alaluf - The CEO of Wand AI, who has championed this work and is a believer in its applicability to our company’s interests.
  • Jennifer L. LeSieur, Ameena Amdahl-Mason, Jason Miller, and Steven Allyn Taylor - High School Teachers and Debate Educators who guided Allen Roush in Policy Debate and served as a catalyst for this work.
  • We also acknowledge the assistance of the National Speech and Debate Association (NSDA), the National Debate Coaches Association (NDCA), the National Debate Tournament (NDT), the American Forensic Association (AFA), and the California National Debate Institute (CNDI).
Appendix
We include additional experiments, which could not fit within the page limits of the original paper in this section.
Summarization Performance on xSum
The xSum dataset (Narayan et al., 2018) (Table 5) comprises 11,334 validation documents and is designed for single-sentence news summarization. Unlike the argumentative datasets, xSum focuses solely on factual accuracy and conciseness without requiring support quality. Our experiments on xSum demonstrate that the fine-tuning techniques, particularly LoRA, consistently improve summarization performance across different model sizes.
Table 5: Performance on xSum (Validation Set, 11,334 documents). ROUGE F1 scores and perplexity on the validation set. Scores are averaged over three runs. R-1, R-2, and R-L denote ROUGE-1, ROUGE-2, and ROUGE-L respectively. Error bars represent one standard error over 3 trials.
Model R-1 (%) R-2 (%) R-L (%) Perplexity Output Quality
Mistral-7B
Base $39.2\pm 0.7$ $16.5\pm 0.6$ $31.5\pm 0.8$ $34.7\pm 2.8$ $7.6\pm 0.2$
LoRA $\mathbf{40.8\pm 0.6}$ $\mathbf{17.2\pm 0.5}$ $\mathbf{32.7\pm 0.7}$ $\mathbf{30.5\pm 2.5}$ $7.7\pm 0.2$
ReFT $40.5\pm 0.7$ $17.1\pm 0.5$ $32.5\pm 0.7$ $31.8\pm 2.6$ $7.7\pm 0.2$
LLaMA3-8B
Base $42.0\pm 0.6$ $18.0\pm 0.6$ $34.0\pm 0.8$ $29.2\pm 2.4$ $7.8\pm 0.2$
LoRA $\mathbf{43.7\pm 0.5}$ $\mathbf{19.1\pm 0.4}$ $\mathbf{35.2\pm 0.7}$ $\mathbf{26.8\pm 2.2}$ $8.0\pm 0.2$
ReFT $43.4\pm 0.6$ $19.0\pm 0.5$ $35.0\pm 0.8$ $27.3\pm 2.3$ $8.0\pm 0.2$
LLaMA3-70B
Base $46.5\pm 0.7$ $21.2\pm 0.6$ $38.0\pm 0.8$ $22.4\pm 2.0$ $8.4\pm 0.2$
LoRA $\mathbf{48.2\pm 0.6}$ $\mathbf{22.5\pm 0.5}$ $\mathbf{39.5\pm 0.7}$ $\mathbf{19.6\pm 1.8}$ $\mathbf{8.6\pm 0.2}$
ReFT $47.9\pm 0.7$ $22.3\pm 0.5$ $39.2\pm 0.8$ $20.5\pm 1.9$ $8.5\pm 0.2$
Argument Detection and Classification
We performed evaluations on argument detection and classification tasks using the same fine-tuning techniques—LoRA, ReFT, and Orthogonalization—as in our original experiments. We used evidence from OpenDebateEvidence which randomly sampled. The evidence was labeled based on if the category string was included within the hat, pocket, or label. The objective of these tasks was to classify arguments into one of 4 specified categories: “Topicality,” “Disadvantages,” “Advantages,” and “Counterplans.” The following table summarizes model performance on a training set of $10,000$ examples and a validation set of $1,000$ examples, reporting accuracy, F1 score, precision, and recall. Results are averaged over three runs with one standard error.
Table 6: Model Performance on Argument Detection and Classification Tasks - Validation scores on a training set of 10,000 examples and a validation set of 1,000 examples. Results are averaged over 3 runs with error reported as $\pm$ one standard error.
Model Technique Accuracy (%) F1 Score Precision Recall
Mistral-7B Base $72.8\pm 0.5$ $0.70\pm 0.04$ $0.68\pm 0.03$ $0.69\pm 0.03$
LoRA $\mathbf{78.2\pm 0.4}$ $\mathbf{0.76\pm 0.03}$ $\mathbf{0.75\pm 0.04}$ $\mathbf{0.76\pm 0.04}$
ReFT $77.9\pm 0.6$ $0.75\pm 0.02$ $0.74\pm 0.03$ $0.75\pm 0.03$
Orthogonal $73.1\pm 0.6$ $0.71\pm 0.03$ $0.70\pm 0.03$ $0.71\pm 0.03$
LLaMA3-8B Base $74.5\pm 0.6$ $0.72\pm 0.03$ $0.70\pm 0.02$ $0.71\pm 0.03$
LoRA $\mathbf{81.2\pm 0.3}$ $\mathbf{0.79\pm 0.02}$ $\mathbf{0.78\pm 0.03}$ $\mathbf{0.79\pm 0.02}$
ReFT $80.8\pm 0.5$ $0.78\pm 0.03$ $0.77\pm 0.02$ $0.78\pm 0.03$
Orthogonal $75.2\pm 0.7$ $0.73\pm 0.04$ $0.71\pm 0.03$ $0.72\pm 0.04$
Stance Detection Performance
We expanded our experimentation to assess model performance on stance detection tasks, using a sampled subset of OpenDebateEvidence with $10{,}000$ examples in the training set and $1{,}000$ for validation. This experiment aims to examine each model’s ability to classify arguments as either "affirmative" (aff) or "negative" (neg), focusing on models fine-tuned with different techniques. We label each row based on "side" dataset column.
Results, displayed in Table 7, show that the LLaMA3-8B models consistently outperformed the Mistral-7B models across all metrics, particularly in accuracy and F1-score. Fine-tuning with LoRA and ReFT provided notable improvements over the base models, with LoRA yielding the highest accuracy and F1-scores on LLaMA3-8B when using $1{,}000{,}000$ examples for fine-tuning. This reinforces LoRA’s effectiveness in parameter adaptation and generalization across diverse tasks. ReFT also showed substantial gains in stance detection, suggesting robustness in refining hidden representations for classification tasks. Orthogonalization provided moderate improvements, but its impact was less pronounced than LoRA and ReFT, potentially due to its paramater subtractive technique.
These findings highlight the importance of fine-tuning techniques, especially LoRA and ReFT, in enhancing model performance on stance detection. The LoRA fine-tuned LLaMA3-8B model, using an extended dataset, achieved an accuracy of $87.1\%\pm 0.9$ and an F1-score of $85.6\pm 0.6$, the highest among all models, underscoring the technique’s adaptability and efficiency.
Table 7: Performance of Models on Stance Detection Tasks - 10,000 examples in the training set, 1,000 validation examples, validation scores reported. Outputs were constrained to either "aff" or "neg".
Model Technique Accuracy (%) Precision Recall F1-Score
Mistral-7B Base $77.3\pm 0.6$ $75.8\pm 0.7$ $76.2\pm 0.5$ $76.0\pm 0.4$
Mistral-7B LoRA $82.4\pm 0.5$ $81.2\pm 0.6$ $81.7\pm 0.4$ $81.4\pm 0.5$
Mistral-7B ReFT $81.9\pm 0.4$ $80.7\pm 0.5$ $81.1\pm 0.6$ $80.9\pm 0.5$
Mistral-7B Orthogonal $78.8\pm 0.5$ $77.2\pm 0.6$ $77.5\pm 0.4$ $77.3\pm 0.4$
LLaMA3-8B Base $79.5\pm 0.7$ $77.9\pm 0.8$ $78.2\pm 0.7$ $78.0\pm 0.6$
LLaMA3-8B LoRA $85.3\pm 0.8$ $83.8\pm 0.7$ $84.1\pm 0.5$ $83.9\pm 0.6$
LLaMA3-8B ReFT $84.1\pm 0.6$ $82.5\pm 0.7$ $83.0\pm 0.6$ $82.7\pm 0.5$
LLaMA3-8B Orthogonal $80.5\pm 0.7$ $79.2\pm 0.6$ $79.5\pm 0.6$ $79.3\pm 0.5$
LLaMA3-8B LoRA (1M Ex) $87.1\pm 0.9$ $85.4\pm 0.8$ $85.8\pm 0.7$ $85.6\pm 0.6$
Using OpenDebateEvidence for Pre-Training
We evaluated OpenDebateEvidence pretrained vs non-pretrained versions of the Mistral-7B and LLaMA3-8B models on several downstream tasks, including DebateSum, ArgAnalysis35K, and Multi-LexSum. The results indicate that pretraining contributes significantly to improved performance across ROUGE scores, perplexity (PPL), and LLM as Judge (LJ) scores for both output quality (LJ-Q) and support quality (LJ-S). These findings reinforce the value of pretraining for enhancing model capabilities in argument classification, summarization, and multi-domain summarization tasks.
In Table 8, we report the averaged performance metrics across three runs, using a plus/minus notation to denote the standard error. Notably, pretrained versions of both Mistral-7B and LLaMA3-8B consistently outperform their non-pretrained counterparts across all datasets. The Mistral-7B model showed considerable improvements on the DebateSum and Multi-LexSum tasks when pretrained, achieving substantial gains in R-1, R-2, and R-L scores while reducing perplexity by over $60\%$. Similarly, pretraining LLaMA3-8B on these tasks led to significant gains, especially on the ArgAnalysis35K dataset, with ROUGE scores increasing by more than 4 points across all metrics.
These results underscore the importance of targeted pretraining in enhancing performance on specific tasks. The improvements in support quality (LJ-S) and output quality (LJ-Q) further suggest that pretraining helps the models generate more coherent and contextually relevant summaries, which are essential in tasks requiring argument understanding and summarization.
Table 8: Performance of Pretrained and Non-Pretrained Models on Downstream Tasks - The entire dataset is used for both pretraining and fine-tuning. ArgAnalysis35K is an argument quality classification task.
Model Dataset PT R-1 R-2 R-L PPL LJ-Q LJ-S Steps
Mistral-7B DebateSum No 26.3 $\pm$ 4 7.5 $\pm$ 3 23.1 $\pm$ 6 130.5 $\pm$ 42 7.3 $\pm$ 3 7.0 $\pm$ 3 60k
Mistral-7B DebateSum Yes 30.4 $\pm$ 1.4 9.4 $\pm$ 0.9 26.5 $\pm$ 1.1 19.8 $\pm$ 1.7 7.8 $\pm$ 0.3 7.6 $\pm$ 0.3 40k
LLaMA3-8B DebateSum No 24.2 $\pm$ 5 6.9 $\pm$ 3 21.9 $\pm$ 8 95.7 $\pm$ 4.5 7.0 $\pm$ 3 6.7 $\pm$ 3 60k
LLaMA3-8B DebateSum Yes 32.2 $\pm$ 1.5 9.9 $\pm$ 1.0 27.4 $\pm$ 1.2 21.8 $\pm$ 2.5 8.0 $\pm$ 0.2 7.8 $\pm$ 0.2 35k
Mistral-7B ArgAnalysis35K No 44.8 $\pm$ 0.3 21.2 $\pm$ 0.5 40.5 $\pm$ 0.8 25.2 $\pm$ 1.1 7.2 $\pm$ 0.3 7.0 $\pm$ 0.3 50k
Mistral-7B ArgAnalysis35K Yes 48.2 $\pm$ 1.5 24.9 $\pm$ 1.0 43.4 $\pm$ 1.2 21.8 $\pm$ 0.5 7.8 $\pm$ 0.2 7.6 $\pm$ 0.2 35k
LLaMA3-8B ArgAnalysis35K No 42.4 $\pm$ 6 19.6 $\pm$ 2 38.8 $\pm$ 1.3 27.3 $\pm$ 1.3 7.0 $\pm$ 3 6.8 $\pm$ 3 50k
LLaMA3-8B ArgAnalysis35K Yes 48.9 $\pm$ 1.7 25.1 $\pm$ 1.2 44.2 $\pm$ 1.4 22.7 $\pm$ 0.7 8.0 $\pm$ 0.3 7.8 $\pm$ 0.3 30k
Mistral-7B Multi-LexSum No 40.1 $\pm$ 5 20.3 $\pm$ 4 37.7 $\pm$ 9 30.1 $\pm$ 2.3 7.1 $\pm$ 2 7.0 $\pm$ 2 65k
Mistral-7B Multi-LexSum Yes 45.2 $\pm$ 1.2 22.5 $\pm$ 0.7 40.8 $\pm$ 1.0 22.0 $\pm$ 1.1 7.7 $\pm$ 0.3 7.5 $\pm$ 0.3 45k
LLaMA3-8B Multi-LexSum No 38.2 $\pm$ 8 19.5 $\pm$ 3 36.0 $\pm$ 1.0 32.5 $\pm$ 2.5 7.0 $\pm$ 2 6.8 $\pm$ 2 65k
LLaMA3-8B Multi-LexSum Yes 46.8 $\pm$ 1.4 23.0 $\pm$ 0.8 41.7 $\pm$ 1.3 23.9 $\pm$ 1.4 7.8 $\pm$ 0.3 7.6 $\pm$ 0.3 40k
Inter-Annotator Agreement
To assess the reliability of our evaluation metrics, we measured the agreement rates between human annotators and between humans and GPT-4 in scoring summary quality. Table 9 presents the inter-human correlation and human-GPT-4 correlation for different debate formats. The results indicate a high level of agreement both among human annotators and between humans and GPT-4, validating the consistency and reliability of our evaluation approach.
Table 9: Inter-Annotator Agreement Rates and Human-GPT-4 Agreement Rates (50 samples, 10 individuals total). Average scores reported.
Debate Format Team 1 (Argument Quality) Team 1 (Support Score) Team 2 (Argument Quality) Team 2 (Support Score) Inter-Human Corr. (AQ) Inter-Human Corr. (SQ) Human-GPT-4 Corr. (AQ) Human-GPT-4 Corr. (SQ)
Policy Debate $7.4$ $7.8$ $7.3$ $7.7$ $0.81$ $0.79$ $0.78$ $0.75$
Lincoln-Douglas Debate $8.1$ $8.4$ $8.0$ $8.2$ $0.84$ $0.82$ $0.82$ $0.80$
Inter-Annotator Agreement: Human vs. GPT-4 Ratings on Summaries
In this section, we present the results of our inter-annotator agreement analysis between human expert debaters and GPT-4 on summary evaluations. We measured both Argument Quality (AQ) and Support Quality (SQ) across various debate formats, including Policy, Lincoln-Douglas, and Public Forum. Fifty expert debaters rated six summaries each, with results averaged over three runs. Pearson correlation metrics were calculated to assess agreement between human and GPT-4 ratings.
Table 10: Inter-Annotator Agreement: Human vs. GPT-4 Ratings on Summaries - 50 expert debaters rated 6 summaries each. Scores represent mean values with standard error over three trials.
ID Format Human-AQ GPT4-AQ Human-SQ GPT4-SQ
S1 Policy 8.2 $\pm$ 0.1 8.0 $\pm$ 0.1 8.0 $\pm$ 0.1 7.9 $\pm$ 0.1
S2 Lincoln-Douglas 7.6 $\pm$ 0.1 7.8 $\pm$ 0.1 7.4 $\pm$ 0.1 7.6 $\pm$ 0.1
S3 Public Forum 7.2 $\pm$ 0.1 7.4 $\pm$ 0.1 7.1 $\pm$ 0.1 7.3 $\pm$ 0.1
S4 Policy 8.5 $\pm$ 0.1 8.3 $\pm$ 0.1 8.3 $\pm$ 0.1 8.2 $\pm$ 0.1
S5 Lincoln-Douglas 7.8 $\pm$ 0.1 7.7 $\pm$ 0.1 7.5 $\pm$ 0.1 7.6 $\pm$ 0.1
S6 Public Forum 7.4 $\pm$ 0.1 7.2 $\pm$ 0.1 7.2 $\pm$ 0.1 7.1 $\pm$ 0.1
Table 11: Pearson Correlation Results - Statistical correlation of Argument Quality and Support Quality between human and GPT-4 ratings.
Metric Pearson Correlation
Argument Quality 0.78 $\pm$ 0.02
Support Quality 0.74 $\pm$ 0.02
The results in Table 10 show a strong agreement between human and GPT-4 ratings, with both Argument Quality (AQ) and Support Quality (SQ) ratings closely aligned across different debate formats. The Pearson correlation values presented in Table 11 further confirm this, with high correlation values for both AQ and SQ, indicating that GPT-4’s evaluations closely mirror human assessments.
These results provide further confidence in the reliability of GPT-4’s evaluation metrics when assessing argumentation and support quality, suggesting potential for consistent automated evaluation alongside human judgment in summarization tasks.
Neural Semantic Deduplication
We applied neural semantic deduplication using the gte-base-en-v1.5 model, as recommended by the Massive Text Embedding Benchmark (MTEB) and implemented via the sentence-transformers library [^3] https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5. Rows with a "spoken" summary with a similarity score above 0.95 were identified as near-duplicates and removed. Table 12 summarizes the changes in document count after null removal and deduplication. We note that most "duplicates" still have value since rows with duplicate documents often have very different metadata (for example, a different debater using the same piece of evidence).
Table 12: Neural Semantic Deduplication Results - Summary of document counts and duplicate cluster statistics after deduplication.
Metric Value
Initial number of Rows in OpenDebateEvidence 4,830,561
Initial Documents in OpenDebateEvidence (Valid fulltext column) 3,512,280
Documents after Nulls in other columns Removed 2,634,023
Documents after Deduplication 692,989
Percentage of Documents Removed 77.4%
Total Duplicate Clusters Identified 10,819,328
Average Cluster Size 30
Largest Cluster Size 2,081
Cross-Dataset Deduplication Analysis
To ensure the uniqueness of OpenDebateEvidence within the context of other argumentation and legislative datasets, we performed a cross-dataset deduplication analysis. We identified 31,353 overlapping documents (4.01%) with the DebateSum dataset, primarily comprising Policy Evidence (96%) and Lincoln-Douglas Evidence (4%). No overlap was detected with BillSum, confirming OpenDebateEvidence’s distinctiveness for legislative and argumentation-focused tasks. Table 13 provides a summary of the overlap findings.
Table 13: Cross-Dataset Deduplication Analysis - Overlap between OpenDebateEvidence and other datasets.
Dataset Overlapping Documents Percentage Overlap (%)
DebateSum 31,353 4.01
BillSum 0 0.00
These deduplication efforts, removing nearly 78% of near-duplicate content, alongside cross-dataset comparisons, affirm OpenDebateEvidence’s uniqueness and utility in argumentative tasks, with minimal redundancy or overlap with other prominent datasets. The deduplicated dataset is publicly available for research and can be accessed at https://huggingface.co/datasets/Hellisotherpeople/OpenCaseList-Deduplicated.
@@ A.1 Representation Bias
While the OpenDebateEvidence dataset is extensive, it may not fully represent the diversity of argumentation styles and topics across all debate communities. The dataset primarily includes evidence from American high school and college debates and, therefore, might not capture the nuances of debates in other cultural or educational contexts or in other languages.
@@ Appendix B Ethics Statement
The OpenDebateEvidence dataset presented in this paper derives from openly shared debate evidence across various educational forums and debate formats. This dataset strictly adheres to the principles of fair use, focusing on academic and research intent. The files that make up OpenDebateEvidence have been hosted online in some cases for over a decade without any known ethical issues arising as a result of it.
@@ Appendix C Social Impacts
In summary, we believe our work will accelerate research, improve model evaluation and training, and ultimately enhance the capabilities of language models in handling complex argumentative texts with no foreseeable negative societal impacts.
@@ D.3 Orthogonalization
We used this notebook for performing this process.
@@ E.1.2 LLM as Judge Prompt
Score the abstract from 0-10 on it’s how well it supports the documents argument, and on its general quality.
@@ E.2.2 LLM as Judge Prompt
SYSTEM PROMPT: You are a lawmaker. USER PROMPT:
@@
Score the abstract from 0-10 on how well it supports the documents argument, and on its general quality.
@@ F.1.1 Dataset Summary
  • Massively increased size (about 25X the size of DebateSum), including nearly all debate evidence ever open-sourced over the past 20 years from High School and College Public Forums, Policy, and Lincoln Douglas debate leagues
@@ F.1.2 Supported Tasks and Leaderboards
This dataset is useful for text generation, summarization, information retrieval, question answering, and related tasks. This dataset is further highly useful as a "trustworthy" dataset. All evidence within it has corresponding citations and is, in general, "factual" or grounded in facts. We do the evaluation in our paper, establishing the first "leaderboard" for measuring the performance of models trained on this dataset.
@@ F.1.3 Languages
English with very minor exceptions (i.e., evidence from performance cases using non-English evidence to make anti-colonialist arguments)
@@ F.1.6 Dataset Format
This dataset was originally contained in CSV files, which were auto-converted into the parqueet dataset format by Huggingface. It’s available for download and consumption in both formats.
@@ F.1.8 Discussion of Biases
Competitive debate at the highest levels has increasingly rewarded teams who cite particular subfields of philosophy. A partial list of these highly represented topics is given below.
@@
These cannons are dominated by so-called "left-wing" thinkers and have mostly marginalized so-called "right-wing" thinkers within them with some notable exceptions
@@ F.1.11 Personal Information
tag Biased abstractive summary of the evidence/argument made by the debater with evidence.
@@
pocket String indicating the virtual “pocket” (top-level section, usually the speech name) in which the
@@
hat String indicating the virtual “hat” (medium-level section, usually the broad type of argument)
@@
block String indicating the virtual “block” (low-level section, usually the specific type of argument)
@@
Year (Top 5) 2022 861,774
2020 850,649
2021 787,765
2019 609,766
2018 423,395
Event cx 2,768,419
ld 1,526,383
pf 43,131
Caselist DisplayName (Top 3) HS LD 2020-21 383,524
HS LD 2021-22 381,591
HS LD 2020-21 326,172
TeamSize 2 2,811,550
1 1,526,383
Level hs 2,767,070
college 1,570,863
State (Top 5) None 2,433,648
CA 622,059
TX 507,454
IL 153,293
GA 127,500
Side N (neg) 2,437,000
A (aff) 1,900,933
DuplicateCount (Top 5) 1 1,639,292
2 189,148
3 136,422
4 110,012
5 91,405